OpenStudy (anonymous):

two forces of numerically equal to 10dynes are acting ,then the resultant is

OpenStudy (jack1):

... im sorry could you please explain your question a bit?

OpenStudy (jack1):

i think it's missing something in the following gaps: two forces of ______ numerically equal to 10dynes are acting ___________ ,then the resultant is

OpenStudy (anonymous):

jack my question is that , what is the resultant if two forces equal to 10dynes are acting at right angles ??

OpenStudy (anonymous):

i think first we shuld dyne convert into newton ....

OpenStudy (jack1):

gotcha, it was missing the rightangles bit before simple triangle resultant then: 1; 1; hypotenuse = root 2 so in this case : 

OpenStudy (anonymous):

yes i think that is a correct answer , thanks jack for the help ..

OpenStudy (jack1):

all good dude

OpenStudy (anonymous):

resultant is \[\sqrt{10^{2} + 10^{2}} = \sqrt{200} = 10\sqrt{2} = 14\]

OpenStudy (jack1):

good catch @JOELMATHEWS1234 , my bad

OpenStudy (anonymous):

thanks , bruv

OpenStudy (anonymous):

better luck next time !

OpenStudy (anonymous):

a) 10dynes b) 20dynes c) 10(3)^1/2 dyne d) 5dynes WHICH OF THE FOLLOWING IS RIGHT ??